# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ Appendix D Proofs
Recall that we consider making a change of measure $\frac{{\mathrm{d}}\mathbb{Q}}{{\mathrm{d}}\mathbb{P}}(X)\propto F(X_{T}).$ We first provide a brief lemma concerning conditional expectations under this change of measure, which we will utilise to prove Theorem 2.1.
Lemma D.1.
@@
As we are considering the change of measure $\frac{{\mathrm{d}}{\mathbb{Q}}}{{\mathrm{d}}\mathbb{P}}(X)\propto F(X_{T})$, a quick calculation shows that the local densities are given by $\tfrac{{\mathrm{d}}{\mathbb{Q}}}{{\mathrm{d}}\mathbb{P}}|_{\sigma(X_{t})}=F_{t}$. For a test function $g(X_{t})$, we therefore have
@@
Here, we used properties of conditional expectations in the second and final lines, and in the first and third lines used the definition of the reweighted measure $\mathbb{Q}$}. We thus see that $\tfrac{1}{F_{t}}{\mathbb{E}}_{\mathbb{P}}[FH|X_{t}]={\mathbb{E}}_{{\mathbb{Q}}}[H|X_{t}]$} $\mathbb{Q}$}-almost surely, as required. ∎
For the next theorem, we assume for the first part that $F$ is differentiable. The second part holds for more general $F$, including for $F=\delta_{x_{T}}$ corresponding to diffusion bridges. See 2.1
@@
We obtain the first statement of the theorem by making the following calculations, starting from Doob’s $h$-transform and using Lemma D.1 with $H=\nabla\log F(X_{T})$ in the final line.
@@
Each of these steps are self-explanatory, with the exception of the interchanging of the derivative and the expectation in the third line. This step is nevertheless intuitive and is well-known in stochastic analysis, and is often referred to as the reparameterisation trick in the machine learning literature. It can be made rigorous using the theory of stochastic flows [25, 20]; namely, if one considers a fixed realisation $\omega$ of the Brownian path between $t$ and $T$, then the terminal state $X_{T}$ is a function of the value $x$ at time $t$, which we can denote $x\mapsto\varphi_{t,T}(x,\omega)=X_{T}(\omega).$ Under mild assumptions on $b,\sigma$ (see [26]), the stochastic flow $\varphi_{t,T}(\cdot,\omega)$ is continuously differentiable. Using Leibniz’s rule, for continuously differentiable $F$ we then have
@@
$=$ ${\mathbb{E}}_{\mathbb{P}}[J_{T|t}^{\top}\,\nabla F(X_{T})|X_{t}=x]$ (55)
@@
[…]
D.1 Proof of Theorem 3.1
In this section we will prove a slightly stronger result, from which Theorem 3.1 will follow. We first start with a definition.
@@
  • bridge-preserving, if the bridges of (3) coincide with those of (1),
  • self-consistent, if (6) is satisfied,
@@
$\mathcal{L}=\sum_{i}b_{i}\partial_{i}+\tfrac{1}{2}\sum_{ij}(\sigma\sigma^{\top})_{ij}\partial_{ij}.$ (61)
@@
    the vector field $u$ is of gradient form and self-consistent.
@@
(i) $\iff$ (ii): This equivalence is well known (and SDEs with bridge preserving drifts are in the reciprocal class of the base process), see for example [28, 38], or [43, Section E.5 ].
@@
On the other hand, we can compute the time evolution of $t\mapsto J_{t|s}^{\top}\,u(t,X_{t}^{u})$ along the controlled dynamics, using the product rule of (Itô) stochastic calculus:
${\mathrm{d}}$ $(J_{t|s}^{\top}\,u(t,X^{u}_{t}))=J_{t|s}^{\top}\,\,{\mathrm{d}}u(t,X^{u}_{t})+{\mathrm{d}}J_{t|s}^{\top}\,u(t,X^{u}_{t})\,$ (62a)
@@
$\qquad+J_{t|s}^{\top}\nabla b(t,X_{t}^{u})\cdot u(t,X_{t}^{u})\,{\mathrm{d}}t$ (62b)
@@
The last line of this calculation uses $\tfrac{1}{2}\nabla|u|^{2}=(u\cdot\nabla)u$, which holds for vector fields of gradient form. Note that in the general case with a possibly spatially-dependent diffusion coefficient, performing similar calculations gives the expression
[…]
$+J_{t|s}^{\top}\,(\nabla u\,\sigma+u\,\nabla\sigma)(t,X_{t}^{u})\cdot{\mathrm{d}}B_{t}.$ (63)
The equivalence between (ii) and (iii) essentially follows from the observation that the drift in ${\mathrm{d}}(J_{t|s}^{\top}\,u(t,X^{u}_{t}))$ coincides with the right-hand side of (Grad-HJB). More precisely, from the above calculation the self-consistency property (6) can be re-expressed in the form
$u(s,x)$ $={\mathbb{E}}_{\mathbb{P}^{u}}[J_{t|s}^{\top}\,u(t,X_{t})|X_{s}=x]$ (64)
$={\mathbb{E}}_{\mathbb{P}^{u}}\left[J_{s|s}^{\top}\,u(s,X^{u}_{s})+\int_{s}^{t}\left(\partial_{t}u+\nabla(b\cdot u)+\tfrac{\sigma^{2}}{2}\Delta u+\tfrac{\sigma^{2}}{2}\nabla|u|^{2}\right)(r,X_{r}^{u})\,{\mathrm{d}}r\Big|X^{u}_{s}=x\right],$ (65)
@@
${\mathbb{E}}_{\mathbb{P}^{u}}\left[\int_{s}^{t}\left(\partial_{t}u+\nabla(b\cdot u)+\tfrac{\sigma^{2}}{2}\Delta u+\tfrac{\sigma^{2}}{2}\nabla|u|^{2}\right)(r,X_{r}^{u})\,{\mathrm{d}}r\Big|X^{u}_{s}=x\right]=0.$ (66)
@@
$\left(\partial_{t}+(b+u)\cdot\nabla+\tfrac{1}{2}\Delta\right)\Omega+\left(\nabla(b+u)\right)\Omega+\Omega\left(\nabla(b+u)\right)^{\top}=0,$ (69)
which clearly admits $\Omega\equiv 0$ as a solution. Under standard regularity assumptions, solutions to (69) are unique once $\Omega(t,\cdot)$ is prescribed for some $t$ (both forwards and backwards in time), see for example [33]. From this, it follows that (iv) implies (iii), and the converse implication follows directly from the definition. ∎
@@
D.2 Generalised self-consistency property
In this section, we provide the proofs of the analogous results for the generalised self-consistency property (20) presented in Appendix B.4. We first fill in the details of the change-of-measure argument outlined in Appendix B.4. Recall that $\tilde{J}_{t|s}$ is the Jacobian process associated to the auxiliary process $\tilde{X}$, which evolves according to the SDE ${\mathrm{d}}\tilde{X}_{t}=(b+\tilde{b})(t,\tilde{X}_{t}){\mathrm{d}}t+\sigma(t,\tilde{X}_{t}){\mathrm{d}}B_{t}$ that includes an additional drift function $\tilde{b}$.
See B.1
$u^{*}(s,x)={\mathbb{E}}_{\mathbb{P}^{u^{*}}}\Big[-\int_{s}^{t}$ $\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+\tilde{J}_{t|s}^{\top}\,u(t,X_{t})|X_{s}=x\Big].$ (70)

### v2

@@ B.2 Solving for the training targets
  • Average prediction: This choice instead aims to enforce the self-consistency uniformly across the remaining time $t\in[s,T]$, so $\mathcal{U}_{s}=\tfrac{1}{T-s}\int_{s}^{T}J_{t|s}^{\top}\,u(t,X_{t}^{u}){\mathrm{d}}t$. In this case, the recursive formula becomes
    $\mathcal{U}_{s}=J_{s+\delta t|s}^{\top}\,\big[\tfrac{T-(s+\delta t)}{T-s}\mathcal{U}_{s+\delta t}+\tfrac{(\delta t)}{T-s}u(s+\delta t,X_{s+\delta t}^{u})\big].$ (24)
    We found this choice to provide a good balance between training target variance and propagation of the terminal information, and it worked consistently well across the experimental settings (see results in Section E.7). For this reason, the results reported in the main body use the average $\alpha$-schedule.
  • Endpoint prediction: This choice chooses to only enforce the self-consistency property to the terminal step; that is, it tries to enforce $u(s,X_{s}^{u})=\mathbb{E}[J_{T-\delta t|s}^{\top}\,u(T-\delta t,X_{T-\delta t}^{u})|X_{s}^{u}]$. This corresponds to taking $\lambda_{s}=1$ in (22), so the targets are calculated as
    $\mathcal{U}_{s}=J_{s+\delta t|s}^{\top}\,\mathcal{U}_{s+\delta t}.$ (25)
    While this choice focuses on using the known terminal information, variance can accrue in the many intermediate steps. Moreover, target variance is exacerbated in the case of diffusion bridges, as the control grows sharply towards the terminal time. In the experiments in Section E.7, we consistently found this choice to perform worse than average prediction.
The above recursions can be understood as a backwards discretisation of continuous dynamics. If we consider the integral $I_{s}=\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}u(t,X_{t}^{u})\;{\mathrm{d}}t$, then this has the continuous dynamics ${\mathrm{d}}I_{s}=-\alpha_{s}\,u(s,X_{s}^{u})\,{\mathrm{d}}s-\nabla b(s,X_{s}^{u})^{\top}I_{s}\,{\mathrm{d}}s$, for $\sigma$ not dependent on $x$ (with an additional term if $\sigma$ has $x$-dependence).
B.3 Sticking-the-Landing adjustments
Here, we explain how ‘Sticking-the-Landing’ (STL) adjustments (Roeder et al., 2017) can be included in the training target recursions, similar to the adjustments discussed in Domingo-Enrich (2024). The idea of the STL adjustments is to remove the variance in the training targets at the solution, by leveraging the knowledge that the corresponding potential must solve the appropriate HJB equation. Below, we provide a sketch proof showing how the STL adjustments arise in our framework, and show how they can be incorporated into our algorithm. For full details, see the proof in Appendix D. Note that this presentation is for a scalar noise term $\sigma$; the general case follows similarly by making the appropriate changes.
It is known that the control $u^{*}$ of the solution is of the form $u^{*}=\nabla\phi$, where $\phi$ solves the HJB equation,
[…]
By taking expectations, we recover the standard self-consistency property (SC1) used in the standard version of our algorithm. However, numerically speaking, working with the expectation introduces variance, as the stochastic term must be averaged out across samples. Rather than working with the standard self-consistency property, ‘sticking-the-landing’ means instead working with this path-wise property directly, which at optimality will be satisfied almost surely.
Computing the STL training targets
$\mathcal{U}_{s}:=\frac{1}{A_{s}}\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}^{u})\,\mathrm{d}t-\frac{1}{A_{s}}\int_{s}^{T}\alpha_{t}\left(\sigma\int_{s}^{t}J_{r|s}^{\top}\nabla u(r,X_{r}^{u})\cdot\mathrm{d}B_{r}\right)\mathrm{d}t.$ (28)
$\mathcal{U}_{s}$ $=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{\sigma}{A_{s}}\int_{s}^{T}A_{r}J_{r|s}^{\top}\nabla u(r,X_{r})\cdot\mathrm{d}B_{r}$ (29)
$=\Big[\tfrac{1}{A_{s}}\int_{s+\delta t}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{\sigma}{A_{s}}\int_{s+\delta t}^{T}A_{r}J_{r|s}^{\top}\nabla u(r,X_{r})\cdot\mathrm{d}B_{r}\Big]$ (30)
$\qquad\qquad+\Big[\tfrac{1}{A_{s}}\int_{s}^{s+\delta t}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{\sigma}{A_{s}}\int_{s}^{s+\delta t}A_{r}J_{r|s}^{\top}\nabla u(r,X_{r})\cdot\mathrm{d}B_{r}\Big]$ (31)
$=J_{s+\delta t|s}^{\top}\,\Big[\tfrac{1}{A_{s}}\int_{s+\delta t}^{T}\alpha_{t}J_{t|s+\delta t}^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{\sigma}{A_{s}}\int_{s+\delta t}^{T}A_{r}J_{t|s+\delta t}^{\top}\nabla u(r,X_{r})\cdot\mathrm{d}B_{r}\Big]$ (32)
$\qquad\qquad+\Big[\tfrac{1}{A_{s}}\int_{s}^{s+\delta t}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{\sigma}{A_{s}}\int_{s}^{s+\delta t}A_{r}J_{r|s}^{\top}\nabla u(r,X_{r})\cdot\mathrm{d}B_{r}\Big]$ (33)
$\approx\sigma\nabla u(s,X_{s})\cdot\delta B_{s}+J_{s+\delta t|s}^{\top}\,\big[\tfrac{A_{s+\delta t}}{A_{s}}\mathcal{U}_{s+\delta t}+(\delta t)\tfrac{\alpha_{s+\delta t}}{A_{s}}u(s+\delta t,X_{s+\delta t})\big].$ (34)
The only change to the standard case is the inclusion of the additional STL adjustment $\sigma\nabla u(s,X_{s})\cdot\delta B_{s}$ in the recursion, which is consistent with the calculations in Domingo-Enrich (2024). In the case of a possibly spatially-dependent diffusion coefficient $\sigma$, performing similar calculations gives the path-wise self-consistency property
Performance using STL adjustments
We compare performance with and without the STL adjustments in Section E.8, and observe that it generally improves performance by a small amount, with the largest improvements seen in cases where large Jacobian terms arise (that is, the double-well and Müller-Brown experiments). The additional terms $(\nabla u\,\sigma+u\,\nabla\sigma)(r,X_{r}^{u})\cdot\delta B_{r}$} can be computed efficiently using Jacobian-vector products, however this still adds an additional computational cost; generally we found that including the STL adjustments made the training steps approximately 1.5 times slower.
Appendix C The Generalised Self-Consistency Property and Algorithm
Algorithm 2 Control Consistency Diffusion Bridge (Generalised version)
for $n=0$ to N1N-1 do
  Solve for training targets $\mathcal{U}_{s}$ (41) backwards along the obtained paths $(X_{t}^{\theta})$, using Algorithm 3.
  Perform a gradient step on $\theta$ using the squared loss function
$\mathcal{L}(\theta)=\mathop{{\mathbb{E}}}\limits_{s\sim\mathbf{U}_{[0,T]}}\left[\left\|u_{\theta}(s,X_{s}^{\bar{\theta}})-\mathcal{U}_{s}\right\|^{2}\right].$ (37)
end for
do: Initialise at known final control
$\mathcal{U}_{T-\delta t}=u(T-\delta t,X_{T-\delta t})=(\sigma\sigma^{\top})(T-\delta t,X_{T-\delta t})^{-1}\Big(\frac{X_{T}-X_{T-\delta t}}{\delta t}-b(T-\delta t,X_{T-\delta t})\Big)$ (38)
for $s=T-2\delta t,T-3\delta t,...,\delta t,0$ do
  Compute next training target $\mathcal{U}_{s}$ as
$\mathcal{U}_{s}$ $=-(\delta t)\nabla\tilde{b}(s,X_{s})^{\top}u(s,X_{s})+\tilde{J}_{s+\delta t|s}^{\top}\,\bigg[\tfrac{(\delta t)\alpha_{s+\delta t}}{A_{s}}u(s+\delta t,X_{s+\delta t})+\tfrac{A_{s+\delta t}}{A_{s}}\mathcal{U}_{s+\delta t}\bigg]$ (39)
$\qquad\qquad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+(\nabla u\,\sigma+u\,\nabla\sigma)(s,X_{s}^{u})\cdot\delta B_{t}}$ (Optional STL adjustment)
end for
In this section, we provide more details regarding the generalised form of the self-consistency property outlined in Section 4. Recall we can perform a change of measure
where $\tilde{\mathbb{P}}$ is associated to an ‘auxiliary’ process ${\mathrm{d}}\tilde{X}_{t}=(b+\tilde{b})(t,\tilde{X}_{t}){\mathrm{d}}t+\sigma(t,\tilde{X}_{t}){\mathrm{d}}B_{t}$, in which we include an additional drift term $\tilde{b}$. The Radon-Nikodym derivative is given by the Girsanov weights $\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}=\exp(-\int_{s}^{T}\sigma^{-1}\tilde{b}(r,X_{r})^{\top}{\mathrm{d}}B_{r}-\int_{s}^{T}\tfrac{(\sigma\sigma^{\top})^{-1}}{2}\|\tilde{b}(r,X_{r})\|^{2}{\mathrm{d}}r)$ (Øksendal, 2003, Theorem 8.6.8). Continuing the calculations from Section 2 (see Section D.3 for a full derivation), we obtain the following generalised self-consistency property.
Theorem C.1 (Generalised self-consistency property).
$u^{*}(s,x)={\mathbb{E}}\Big[-\int_{s}^{t}$ $\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r}^{u^{*}})^{\top}u^{*}(r,X_{r}^{u^{*}}){\mathrm{d}}r+\tilde{J}_{t|s}^{\top}\,u(t,X_{t}^{u^{*}})|X_{s}^{u^{*}}=x\Big],\textrm{~~~~~for~~}0<s<t<T.$ (40)
When $\sigma$ is independent of $x$, taking $\tilde{b}=-b$ yields the simplified form (4.1) presented in the main paper. Theorem 4.2 states that we can instead use (40) in place of the self-consistency property (SC1) in our proposed method, while retaining the same theoretical guarantees. We now provide details for implementing this generalised version of our algorithm.
Solving for the training targets
Unlike in the standard case, the generalised self-consistency property includes ‘running costs’. As before, we can construct the targets by weighting according to an $\alpha$-schedule. Denoting the training targets as
$\mathcal{U}_{s}=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}\big(-\int_{s}^{t}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u(r,X_{r}){\mathrm{d}}r+\tilde{J}_{t|s}^{\top}\,u(t,X_{t})\big){\mathrm{d}}t,$ (41)
we can modify the reverse recursions for computing $\mathcal{U}_{s}$. For completeness, we include the computations below.
$\mathcal{U}_{s}$ $=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}\big(-\int_{s}^{t}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u(r,X_{r}){\mathrm{d}}r+\tilde{J}_{t|s}^{\top}\,u(t,X_{t})\big){\mathrm{d}}t$ (42)
$\qquad\qquad\qquad\qquad+\tfrac{1}{A_{s}}\tilde{J}_{s+\delta t|s}^{\top}\,\int_{s+\delta t}^{T}\alpha_{t}\big(-\int_{s+\delta t}^{t}\tilde{J}_{r|s+\delta t}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u(r,X_{r}){\mathrm{d}}r+\tilde{J}_{t|s+\delta t}^{\top}\,u(t,X_{t})\big){\mathrm{d}}t$ (47)
$\approx-(\delta t)\nabla\tilde{b}(s,X_{s})^{\top}u(s,X_{s})+\tilde{J}_{s+\delta t|s}^{\top}\,\bigg[\tfrac{(\delta t)\alpha_{s+\delta t}}{A_{s}}u(s+\delta t,X_{s+\delta t})+\tfrac{A_{s+\delta t}}{A_{s}}\mathcal{U}_{s+\delta t}\bigg]$ (48)
When using the second form of the self-consistency property, (4.1), this simplifies to
$\mathcal{U}_{s}=(\delta t)\nabla b(s,X_{s}^{u})^{\top}u(s,X_{s}^{u})+\bigg[\tfrac{(\delta t)\alpha_{s+\delta t}}{A_{s}}u(s+\delta t,X_{s+\delta t}^{u})+\tfrac{A_{s+\delta t}}{A_{s}}\mathcal{U}_{s+\delta t}\bigg]$ (49)
Algorithm
A description of the resulting generalised version of the algorithm is given in Algorithm 2. The method for solving for the training targets $\mathcal{U}_{s}$ backwards along the trajectories in the generalised case is given in Algorithm 3. In the paper, we focus on the two most natural forms of the self-consistency property, namely (SC1) and (4.1), which we anticipate will be suitable for the vast majority of applications.
Sticking-the-Landing Adjustments
By examining the proof of Theorem C.1 in Appendix D, we see that the STL adjustments for the generalised version of the algorithm remain the same as the standard case discussed in Section B.3. They are included in blue in Algorithm 3.
@@ Appendix D Proofs
D.1 Details of Sketch Derivation in Section 2
In Section 2, we provided an introductory derivation of the self-consistency property (SC1); here, we fill in the technical details of the calculations. We remark that (SC1) can also be derived directly from the computations in the proof of Theorem 4.2 provided in the next section.
Recall that we consider making a change of measure $\frac{{\mathrm{d}}\mathbb{Q}}{{\mathrm{d}}\mathbb{P}}(X)\propto F(X_{T}).$ We first provide a brief lemma concerning conditional expectations under this change of measure (see Solé (2005, Lemma 4.4)), which is utilised in the sketch derivation in Section 2 and in the full proof of Theorem 2.1.
Lemma D.1 (Conditional change of measure).
@@ D.1 Details of Sketch Derivation in Section 2
Let use consider the change of measure $\frac{{\mathrm{d}}{\mathbb{Q}}}{{\mathrm{d}}\mathbb{P}}(X)=F(X_{T})$ (where we assume the normalising constant is included in $F$; otherwise, one can rescale accordingly). A quick calculation shows that the local densities are given by $\tfrac{{\mathrm{d}}{\mathbb{Q}}}{{\mathrm{d}}\mathbb{P}}|_{\sigma(X_{t})}=F_{t}$. For a test function $g(X_{t})$, we therefore have
@@
Here, we used the tower property of conditional expectations in the second and final lines, and in the first and third lines used the definition of the reweighted measure $\mathbb{Q}$}. Since $g$ is an arbitrary test function, we thus see that $\tfrac{1}{F_{t}}{\mathbb{E}}_{\mathbb{P}}[FH|X_{t}]={\mathbb{E}}_{{\mathbb{Q}}}[H|X_{t}]$} $\mathbb{Q}$}-almost surely, as required. ∎
See 2.1
@@
We begin by filling in the details of the sketch derivation in Section 2 showing that $u^{*}(t,x)={\mathbb{E}}_{\mathbb{Q}}[J_{T|t}^{\top}\,\nabla\log F(X_{T})|X_{t}=x]$ when $F$ is differentiable. Starting from Doob’s $h$-transform and using Lemma D.1 with $H=\nabla\log F(X_{T})$ in the final line, we have
@@
The interchanging of the derivative and the expectation in the third line is often referred to as the reparameterisation trick in the machine learning literature. It can be made rigorous using the theory of stochastic flows (Kunita, 1984; Elworthy et al., 1999); namely, if one considers a fixed realisation $\omega$ of the Brownian path between $t$ and $T$, then the terminal state $X_{T}$ is a function of the value $x$ at time $t$, which we can denote $x\mapsto\varphi_{t,T}(x,\omega)=X_{T}(\omega).$ Under mild assumptions on $b,\sigma$ (see Kunita (1986)), the stochastic flow $\varphi_{t,T}(\cdot,\omega)$ is continuously differentiable. Using Leibniz’s rule, for continuously differentiable $F$ we then have,
@@
$=$ ${\mathbb{E}}_{\mathbb{P}}[J_{T|t}^{\top}\,\nabla F(X_{T})|X_{t}=x].$ (62)
@@
Finally, we remark that this self-consistency property also holds in the singular case when $F=\delta_{x_{T}}$, which corresponds to diffusion bridges. To see this, take $\tilde{T}<T$, and apply the above result with $\tilde{T}$ in place of $T$ and with ${F(X_{\tilde{T}})=p(X_{T}=x_{T}|X_{\tilde{T}})}$, which is smooth. We thus obtain the result for $s<t<\tilde{T}$, and as $\tilde{T}$ was arbitrary we see that the self-consistency property holds for $s<t<T$. ∎
D.2 Proof of Theorem 3.1
We now move to proving the characterisation of the diffusion bridge given in Theorem 3.1, upon which our approach is based. In fact, all results in the paper essentially follow from such computations, which involve relating the self-consistency property to the gradient of the HJB equation by considering an Itô expansion of $J_{t|s}^{\top}\,u(t,X_{t}^{u})$ along the controlled dynamics.
We will prove a slightly stronger result in Theorem D.3, from which Theorem 3.1 will follow. To begin, we introduce the following definitions.
@@
  • bridge-preserving, if the bridges of the controlled dynamics $X_{t}^{u}$ (3) coincide with those of the base dynamics $X_{t}$(1),
  • self-consistent, if $u$ satisfies $u(s,x)={\mathbb{E}}[J_{t|s}^{\top}\,u(t,X_{t}^{u})|X_{s}^{u}=x]$,
@@
$\mathcal{L}=\sum_{i}b_{i}\partial_{i}+\tfrac{1}{2}\sum_{ij}(\sigma\sigma^{\top})_{ij}\partial_{ij}.$
@@
    the vector field $u$ is of gradient form and self-consistent,
@@
(i) $\iff$ (ii): This equivalence is well known (and SDEs with bridge preserving drifts are in the reciprocal class of the base process), see for example Léonard (2013); Rœlly (2013), or Vargas et al. (2024, Section E.5).
@@
On the other hand, we can compute the time evolution of $t\mapsto J_{t|s}^{\top}\,u(t,X_{t}^{u})$ along the controlled dynamics, using the product rule of Itô stochastic calculus:
${\mathrm{d}}(J_{t|s}^{\top}\,u(t,$ $X^{u}_{t}))=J_{t|s}^{\top}\,\,{\mathrm{d}}u(t,X^{u}_{t})+{\mathrm{d}}J_{t|s}^{\top}\,u(t,X^{u}_{t})\,$
@@
$\qquad+J_{t|s}^{\top}\nabla b(t,X_{t}^{u})\cdot u(t,X_{t}^{u})\,{\mathrm{d}}t$
@@
The last line of this calculation uses $\tfrac{1}{2}\nabla|u|^{2}=(u\cdot\nabla)u$, which holds for vector fields of gradient form. Note that in the general case with a possibly spatially-dependent diffusion coefficient, performing similar calculations instead gives rise to the noise term $J_{t|s}^{\top}\,(\nabla u\,\sigma+u\,\nabla\sigma)(t,X_{t}^{u})\cdot{\mathrm{d}}B_{t}$.
The equivalence between (ii) and (iii) follows from the observation that the drift in ${\mathrm{d}}(J_{t|s}^{\top}\,u(t,X^{u}_{t}))$ coincides with the right-hand side of (Grad-HJB). More precisely, from the above calculation the self-consistency property (SC1) can be re-expressed in the form
$u(s,x)$ $={\mathbb{E}}[J_{t|s}^{\top}\,u(t,X_{t}^{u})|X_{s}^{u}=x]$
$={\mathbb{E}}\left[J_{s|s}^{\top}\,u(s,X^{u}_{s})+\int_{s}^{t}J_{t|s}^{\top}\,\left(\partial_{t}u+\nabla(b\cdot u)+\tfrac{\sigma^{2}}{2}\Delta u+\tfrac{\sigma^{2}}{2}\nabla|u|^{2}\right)(r,X_{r}^{u})\,{\mathrm{d}}r\Big|X^{u}_{s}=x\right],$
@@
${\mathbb{E}}\left[\int_{s}^{t}\left(\partial_{t}u+\nabla(b\cdot u)+\tfrac{\sigma^{2}}{2}\Delta u+\tfrac{\sigma^{2}}{2}\nabla|u|^{2}\right)(r,X_{r}^{u})\,{\mathrm{d}}r\Big|X^{u}_{s}=x\right]=0.$ (69)
@@
$\left(\partial_{t}+(b+u)\cdot\nabla+\tfrac{\sigma^{2}}{2}\Delta\right)\Omega+\left(\nabla(b+u)\right)\Omega+\Omega\left(\nabla(b+u)\right)^{\top}=0,$ (72)
which clearly admits $\Omega\equiv 0$ as a solution. Under standard regularity assumptions, solutions to (72) are unique once $\Omega(t,\cdot)$ is prescribed for some $t$ (both forwards and backwards in time), see for example (Pazy, 2012). From this, it follows that (iv) implies (iii), and the converse implication follows directly from the definition. ∎
@@
D.3 Generalisations of the Self-Consistency Property
We now provide the proofs of the analogous results for the generalised self-consistency property (40) presented in Section 4 and Appendix C. Recall that $\tilde{J}_{t|s}$ is the Jacobian process associated to the auxiliary process $\tilde{X}$, which evolves according to the SDE ${\mathrm{d}}\tilde{X}_{t}=(b+\tilde{b})(t,\tilde{X}_{t}){\mathrm{d}}t+\sigma(t,\tilde{X}_{t}){\mathrm{d}}B_{t}$ including an additional drift function $\tilde{b}$. Theorems C.1 and 4.2 follow from a minor modification to the proof of Theorem D.3.
See C.1
See 4.2